SIR PLESE SOLVE THIS





THREE LIQUIDS P,Q,R ARE GIVEN. 4KG OF P AT 60 DEGREE C AND 1KG OF R AT 50 DEGREE C WHEN MIXED PRODUCE RESULTANT TEMPRATURE 55 DEGEE C.A MIXTURE OF 1KG OF P AT 60 DEGREE C AND 1KG OF Q AT 50 DEGREE C SHOWS TEMPERATURE 55 DEGREE C.WHAT WILL BE THE RESULTING TEMPERATURE WHEN 1KG OF Q AT 60 DEGEE C IS MIXED WITH 1KG OF R AT 50 DEGREE C?




Grade:11

1 Answers

Rohith Gandhi
14 years ago

 

Dear Rohan,

case 1

4kg of P at 60 degree and 1 kg of R at 50 degree, resultant temp of mix = 55 degree

so, energy lost by P = energy gained by R and from calorimetry,

4XSpX(60-55)=1XSrX(55-50)   =>Sr = 4Sp

case 2

1kg of P at 60 degree and 1 kg of Q at 50 degree, resultant temp of mix = 55 degree

so, energy lost by P = energy gained by Q and from calorimetry,

1XSpX(60-55)=1XSqX(55-50)   =>Sq = Sp

from both above cases, we get, Sr = 4Sq

case 3

1kg of Q at 60 degree and 1 kg of R at 50 degree, resultant temp of mix = ?, let it be T

so, energy lost by Q = energy gained by R and from calorimetry,

1XSqX(60-T)=1XSrX(T-50)   => SqX(60-T)=4SqX(T-50) =>60-T = 4T - 200 =>5T = 260 => T = 260/5 = 52 degree

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Rohan !!!




Regards,

Askiitians Experts

Rohith Gandhi